OpenStudy (anonymous):

how do you solve this? square root of x +8=9

OpenStudy (turingtest):

\[\sqrt x+8=9\]yes?

OpenStudy (turingtest):

or\[\sqrt{x+8}=9\]which one?

OpenStudy (anonymous):



OpenStudy (turingtest):

same as before first get all numbers on the same side of the = sign

OpenStudy (turingtest):

so what should we subtract from both sides?

OpenStudy (anonymous):

the 8 right?

OpenStudy (turingtest):

yes, and what does that give you ?

OpenStudy (anonymous):

got it:)

OpenStudy (turingtest):

excellent :)

OpenStudy (anonymous):

thank you!!! :)

OpenStudy (anonymous):

can u help me solve something else please??